Chivon invests $2757 in a simple interest account. The annual interest rate is 2.1%.


How much money in total will she have earned at the end of 3 years?



1
Expert's answer
2021-04-15T07:25:42-0400

"interest, I=\frac{PRT}{100}=\frac{(2757\times 2.1\times 3)}{100}=\$173.691"

"Amount, A=P+I=2757+173.691=\$2,930.691"